OpenStudy (anonymous):

Which two points satisfy y = -x2 + 2x + 4 and x + y = 4?

OpenStudy (zpupster):

lets graph it first where they intersect is where our 2 points will satisfy

OpenStudy (zpupster):

let x = 0 y= ? here it satisfys one point (x,y) or (0,?)

OpenStudy (zpupster):

x + y = 4 4-x = y subtracting x both sides 4-x = - x2 + 2x + 4 ----substituting 4-x for y -x = -x2 + 2x subtracting 4 both sides 0 = -x2 + 3x adding x 0 = -x(x-3) factoring -x out let both equal zero by Zero Product Property x-3=0 (x,y) so (3,?) substitute back in we already have our zero for x from before